FriskIsNice:

Help me! I’ve been doing this for a while, I just need the coordinates for once, please help! I never understand diagonal lines (will send it now)

FriskIsNice:

surjithayer:

you want the co-ordinates after reflection about blue line.

surjithayer:

or reflection about x-axis or y-axis.

mdobbs6856:

Are you just figuring out what the points are labeled?

surjithayer:

i think you want reflection about blue line. B(-5,0),E(-4,0),C(-5,8),D(-4,8) Eq. of blue line is y=x ...(1) or y-x=0 slope of line y=x is 1. slope of line perpendicular to blue line =-1 eq. of line through (-5,0) with slope -1 is y-0=-1(x+5) y+x+5=0 ...(2) point of intersection of (1) and (2) is x+x+5=0 2x=-5 x=-5/2 y=x=-5/2 let (x,y) be co-ordinates of B'(the reflection of B about blue line) \[\frac{ x+(-5) }{ 2 }=-\frac{ 5 }{ 2 }\] x-5=-5 x=0 \[\frac{ y+0 }{ 2 }=-\frac{ 5 }{ 2}\] y=-5 B'(0,-5) proceeding this way you can find C',D',E'